### v1

# Auslander-Reiten translations in the monomorphism categories of exact categories

Abstract.
Let $\Lambda$ be a finite dimensional algebra. Let ${\mathcal{C}}$ be a functorially finite exact subcategory of $\Lambda$-mod with enough projective and injective objects and ${\mathcal{S}}({\mathcal{C}})$ be its monomorphism category. It turns out that the category ${\mathcal{S}}({\mathcal{C}})$ has almost split sequences. We show an explicit formula for the Auslander-Reiten translation in ${\mathcal{S}}({\mathcal{C}})$. Furthermore, if ${\mathcal{C}}$ is a stably $d$-Calabi-Yau Frobenius category, we calculate objects under powers of Auslander-Reiten translation in the triangulated category $\overline{{\mathcal{S}}({\mathcal{C}})}$.
Key words and phrases.  (mono)morphism categories,  exact categories,  Auslander-Reiten translation,  Auslander-Reiten-Serre duality,  functor categories.
MSC2020: 16G70, 18A20, 18A25, 16B50.
The corresponding author.
Supported by the NSF of China (No. 12201321).
[…]
School of Mathematics and Statistics,   Nantong University
Jiangsu 226019, P. R. China
1. Introduction
Throughout the paper, $\Lambda$ stands for a finite dimensional algebra over a field $k$ and $\Lambda\textendash\operatorname{mod}$ the category of finitely generated left $\Lambda$-modules. Denote by $D=\operatorname{Hom}_{k}(-,k)$ the $k$-dual functor.
1.1. Background
The classification of subobjects is a fundamental problem in mathematics. In [Birk], G. Birkhoff proposed the problem of classifying all subgroups of abelian $p$-groups by matrices. This question is related to representations of partial order sets, such as [Ar, Si1, Si2]. C. M. Ringel and M. Schmidmeier [RS1, RS2, RS3] revisited this subject using a modern point of view by studying the representation type and the Auslander-Reiten theory of the submodule category for artin algebras $k[x]/(x^{n})$, for $n>1$. In [RS2], they developed the Auslander-Reiten theory in the submodule category for an artin algebra and showed that the Ausladner-Reiten translation in the submodule category ${\mathcal{S}}(\Lambda)$ can be computed by the Auslander-Reiten translation within the module category.
Theorem 1.1.
Let $\Lambda$ be an artin algebra. Then the submodule category ${\mathcal{S}}(\Lambda)$ has almost split sequences and the Auslander-Reiten translation in ${\mathcal{S}}(\Lambda)$ satisfies
$\tau_{\mathcal{S}}f=\operatorname{\bf Mimo}\tau_{\Lambda}\operatorname{Coker}f$ (1)
where $\operatorname{\bf Mimo}$ denotes the minimal monomorphism.
This paper is devoted to study the morphism categories ${\mathcal{H}}({\mathcal{C}})$, respectively monomorphism categories ${\mathcal{S}}({\mathcal{C}})$ and epimorphism categories ${\mathcal{F}}({\mathcal{C}})$ of exact categories ${\mathcal{C}}$, where the objects are morphism, respectively inflations and deflations in ${\mathcal{C}}$. It turns out that there have been fruitful applications from this point of view. In [LiZ], the authors showed the following:
Theorem 1.2.
For a Gorenstein algebra $\Lambda$, the category of Gorenstein projective modules over the matrix algebra $T_{2}(\Lambda)=\begin{bmatrix}\Lambda&\Lambda\\ 0&\Lambda\end{bmatrix}$ is ${\rm Gproj}(T_{2}(\Lambda))\cong{\mathcal{S}}({\rm Gproj}(\Lambda))$.
In this case, it is well-known that ${\mathcal{S}}({\rm Gproj}(\Lambda))$ is a functorially finite subcategory of $T_{2}(\Lambda)\textendash\operatorname{mod}$, and hence it has almost split sequences. In [LZ3], the authors showed a similar result about the existence of almost split sequences in monomorphism categories.
Theorem 1.3.
If ${\mathcal{C}}$ is contravariantly finite resolving in subcategory in $\Lambda\textendash\operatorname{mod}$, then so is ${\mathcal{S}}({\mathcal{C}})$ in $T_{2}(\Lambda)\textendash\operatorname{mod}$ and hence it has almost split sequences.
1.2. Main results
Motivated by these results above, we will focus on studying the existence of almost split sequences in the monomorphism category of an exact category and their Auslander-Reiten translations in this paper. We will show an explicit formula for the Auslander-Reiten translation in the monomorphism category ${\mathcal{S}}({\mathcal{C}})$ and the epimorphism category ${\mathcal{F}}({\mathcal{C}})$.
Theorem A.
Let $\Lambda$ be a finite dimensional algebra and ${\mathcal{C}}$ be a functorially finite exact subcategory of $\Lambda\textendash\operatorname{mod}$ with enough projective and injective objects. Then the followings hold:
  1. (1)
    The monomorphism category ${\mathcal{S}}({\mathcal{C}})$ and the epimorphism category ${\mathcal{F}}({\mathcal{C}})$ have almost split sequences.
  2. (2)
    Let $\tau_{\mathcal{C}}$ denote the Auslander-Reiten translation of ${\mathcal{C}}$, then the Auslander-Reiten translations of ${\mathcal{S}}({\mathcal{C}})$ and ${\mathcal{F}}({\mathcal{C}})$ are given by:
    $\tau_{\mathcal{S}}(f)\cong\operatorname{\bf Mimo}\tau_{\mathcal{C}}\underline{\operatorname{\bf Cok}(f)},\ \ \tau_{\mathcal{F}}(f)\cong\operatorname{\bf Cok}\operatorname{\bf Mimo}\tau_{\mathcal{C}}(\underline{f}),$
    where $\underline{\operatorname{\bf Cok}(f)}$ and $\underline{f}$ are the images of the maps in the stable category $\underline{{\mathcal{C}}}$.
We need to clarify the notations as well as make some remarks concerning this result.
  1. (i)
    Exact structures and almost split sequences. When ${\mathcal{C}}$ is an exact category, the morphism category ${\mathcal{H}}({\mathcal{C}})$ and its exact subcategories ${\mathcal{S}}({\mathcal{C}})$, ${\mathcal{F}}({\mathcal{C}})$ have a degreewise exact structure inherited from ${\mathcal{C}}$ (see Definition 3.4). The almost split structure in Theorem A is considered with respect to this exact structure.
    Some special almost split sequences in ${\mathcal{H}}({\mathcal{C}})$, ${\mathcal{S}}({\mathcal{C}})$ and ${\mathcal{F}}({\mathcal{C}})$ are classified in Lemma 3.9 and Lemma 6.3 respectively. The others are so-called cw-exact i.e. degreewise split exact (see Definition 3.1).
  2. (ii)
    RSS-equivalence. Let $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C$ be a conflation in an exact category ${\mathcal{C}}$. We denote by $\operatorname{Coker}f=g$ and $\operatorname{Ker}g=f$. In fact, this induces a pair of equivalence functors $\operatorname{\bf Cok}:{\mathcal{S}}({\mathcal{C}})\to{\mathcal{F}}({\mathcal{C}})$ and $\operatorname{\bf Ker}:{\mathcal{F}}({\mathcal{C}})\to{\mathcal{S}}({\mathcal{C}})$, which are called the RSS-equivalences. Furthermore, if ${\mathcal{C}}$ has enough projective objects, they induce functors on the stable categories $\operatorname{\bf Cok}:\underline{{\mathcal{S}}({\mathcal{C}})}\to\underline{{\mathcal{F}}({\mathcal{C}})}$ and $\operatorname{\bf Ker}:\underline{{\mathcal{F}}({\mathcal{C}})}\to\underline{{\mathcal{S}}({\mathcal{C}})}$.
  3. (iii)
    $\operatorname{\bf Mimo}$ and epivalence. The notion of minimal monomorphism was introduced in [RS2]. For any morphism $f:A\to B$ in a module category $\Lambda\textendash\operatorname{mod}$, the minimal monomorphism $\operatorname{\bf Mimo}f=[f,e]^{T}:A\to B\oplus I\operatorname{Ker}f$, where $e$ is an extension of the injective envelope $i:\operatorname{Ker}f\to I\operatorname{Ker}f$. It was shown that $\operatorname{\bf Mimo}f$ is unique up to isomorphism, independent on the choice of $e$ and it is a minimal right ${\mathcal{S}}(\Lambda)$-approximation of $f$. In [BM], the authors pointed out that $\operatorname{\bf Mimo}$ is a functor ${\mathcal{H}}(\Lambda)\to\overline{{\mathcal{S}}(\Lambda)}$. In [LS, GKKP2], it was shown that right ${\mathcal{S}}({\mathcal{C}})$-approximations $\mathfrak{R}f\to f$ give rise to a functor $\mathfrak{R}:{\mathcal{H}}(\Lambda)\to\overline{{\mathcal{S}}(\Lambda)}$. In this paper, we generalize this notion for a Krull-Schmidt exact category ${\mathcal{C}}$ with enough injective objects and denote the image of right ${\mathcal{S}}({\mathcal{C}})$-approximations of $f$ in the stable subcategory $\overline{{\mathcal{S}}({\mathcal{C}})}$ by $\operatorname{\bf Mimo}f$. Hence, $\operatorname{\bf Mimo}:{\mathcal{H}}({\mathcal{C}})\to\overline{{\mathcal{S}}({\mathcal{C}})}$ is a functor.
    Another important functor is the epivalence $\pi:\overline{{\mathcal{S}}({\mathcal{C}})}\to{\mathcal{H}}(\overline{\mathcal{C}})$ (see Definition 2.5, Theorem 5.6), which sends a monomorphism $f$ to its image $\overline{f}$ in the stable category $\overline{\mathcal{C}}$ and $\pi^{\prime}:\underline{{\mathcal{F}}({\mathcal{C}})}\to{\mathcal{H}}(\underline{{\mathcal{C}}})$ which sends an epimorphism $f$ to its image $\underline{f}$ in $\underline{{\mathcal{C}}}$. Since an epivalence induces a bijection of indecomposable objects between categories, we can define a map $\operatorname{\bf Mimo}$ on the objects of ${\mathcal{H}}(\overline{\mathcal{C}})$ which agrees with the inverse of $\pi$ (Remark 5.9).
  4. (iv)
    Functoriality. We clarify the notations in these two Auslander-Reiten formulas. In general, if a Krull-Schmidt exact category ${\mathcal{C}}$ with enough projective and injective objects has almost split sequences, then the Auslander-Reiten translation is a functor $\tau_{\mathcal{C}}:\underline{{\mathcal{C}}}\to\overline{\mathcal{C}}$. However, our formulas for $\tau_{\mathcal{S}}$ and $\tau_{\mathcal{F}}$ only hold for objects $f$ in $\underline{{\mathcal{S}}({\mathcal{C}})}$ or $\underline{{\mathcal{F}}({\mathcal{C}})}$, but not for morphisms. Well-definedness of the formulas can be checked directly from the following compositions:
    ${\tau_{\mathcal{S}}:\underline{{\mathcal{S}}({\mathcal{C}})}}$${\underline{{\mathcal{F}}({\mathcal{C}})}}$${{\mathcal{H}}(\underline{{\mathcal{C}}})}$${{\mathcal{H}}(\overline{\mathcal{C}})}$${\overline{{\mathcal{S}}({\mathcal{C}})}}$$\scriptstyle{\operatorname{\bf Cok}}$$\scriptstyle{\pi^{\prime}}$$\scriptstyle{\tau_{\mathcal{C}}}$$\scriptstyle{\operatorname{\bf Mimo}}$
    ${\tau_{\mathcal{F}}:\underline{{\mathcal{F}}({\mathcal{C}})}}$${{\mathcal{H}}(\underline{{\mathcal{C}}})}$${{\mathcal{H}}(\overline{\mathcal{C}})}$${\overline{{\mathcal{S}}({\mathcal{C}})}}$${\overline{{\mathcal{F}}({\mathcal{C}})}}$$\scriptstyle{\pi^{\prime}}$$\scriptstyle{\tau_{\mathcal{C}}}$$\scriptstyle{\operatorname{\bf Mimo}}$$\scriptstyle{\operatorname{\bf Cok}}$
  5. (v)
    Comparison of the proofs. Although the formulas seem quite similar than those in Theorem 1.1, we would like to mention that we have to choose a different route of proof in this paper. The original proof in [RS2] used techniques of transferring almost split sequences. Since there are functors $\operatorname{\bf Ker}:{\mathcal{H}}(\Lambda)\to{\mathcal{S}}(\Lambda)$ and $\operatorname{\bf Cok}:{\mathcal{H}}(\Lambda)\to{\mathcal{F}}(\Lambda)$ which preserve certain almost split sequences, one can transfer the problem of calculating the Auslander-Reiten translation in ${\mathcal{S}}(\Lambda)$ and ${\mathcal{F}}(\Lambda)$ to the problem of calculating Auslander-Reiten translation in ${\mathcal{H}}(\Lambda)$, which is equivalent to the module category $T_{2}(\Lambda)\textendash\operatorname{mod}$. Then the Auslander-Reiten translation in ${\mathcal{H}}(\Lambda)$ can be computed as $\tau_{\mathcal{H}}=D\operatorname{Tr}$, where $\operatorname{Tr}$ is the transpose functor on $T_{2}(\Lambda)\textendash\operatorname{mod}$. However, one may immediately realize that substituting ${\mathcal{H}}(\Lambda)$ by ${\mathcal{H}}({\mathcal{C}})$ will cause the problem that there no longer exists a “Nakayama functor” or a “transpose functor” on ${\mathcal{H}}({\mathcal{C}})$, since it is not necessarily a module category.
    To solve this problem, we need to transfer almost split sequences through functor categories. The key observation is due to a recent survey by Hafezi and Eshraghi [HE], where they studied the functor $\Theta:{\mathcal{H}}({\mathcal{C}})\to{\mathcal{C}}\textendash\operatorname{mod}$ from the morphism category to the category of finitely presented functors on an additive category ${\mathcal{C}}$, showing that the functor $\Theta$ preserves cw-exact almost split sequences of ${\mathcal{H}}({\mathcal{C}})$ (see Theorem 3.13).
    We briefly explain the strategy of our proof here. First of all, up to some reductions it suffices to prove the formula $\tau_{\mathcal{F}}(f)\cong\operatorname{\bf Cok}\operatorname{\bf Mimo}\tau_{\mathcal{C}}(\underline{f})$ holds for epimorphisms $f$ which is the end term of a cw-exact almost split sequence (see Section 6.3). We use the stable functor ${\mathcal{F}}({\mathcal{C}})\to{\mathcal{H}}(\underline{{\mathcal{C}}})$ ($f\mapsto\underline{f}$) to transfer cw-exact almost split sequences in ${\mathcal{F}}({\mathcal{C}})$ to cw-exact almost split sequences in ${\mathcal{H}}(\underline{{\mathcal{C}}})$ (Lemma 6.5). And then we can use the functor $\Theta:{\mathcal{H}}(\underline{{\mathcal{C}}})\to\underline{{\mathcal{C}}}\textendash\operatorname{mod}$ to transfer cw-exact almost split sequences in ${\mathcal{H}}(\underline{{\mathcal{C}}})$ to almost split sequences in $\underline{{\mathcal{C}}}\textendash\operatorname{mod}$. Finally, the problem boils down to computing the Auslander-Reiten translation on $\underline{{\mathcal{C}}}\textendash\operatorname{mod}$, which has been investigated by [AR74].
  6. (vi)
    Generalizations and relevant results. In this paper, we focus on the “submodule categories” as its importance of being a prototype of the general (separated) monomorphism categories which have been studied intensively in [LZ1, LZ2, ZX]. Meanwhile, further investigations and generalizations are expected. Possible generalizations are two-folded. On one hand, in contrast of our assumption that ${\mathcal{C}}$ is a subcategory of some ambient module category and has enough projective and injective objects, one may consider the Auslander-Reiten theory of ${\mathcal{S}}({\mathcal{C}})$ for a $\operatorname{Hom}$-finite Krull-Schmidt exact category ${\mathcal{C}}$ having almost split sequences (e.g. the category of coherent sheaves ${\rm coh}\mathbb{P}^{1}$). For such categories ${\mathcal{C}}$, it is beyond our knowledge so far to show that ${\mathcal{S}}({\mathcal{C}})$ also has almost split sequences. On the other hand, one may consider the Auslander-Reiten theory for separated monomorphism categories ${\rm smon}(Q,{\mathcal{C}})$ of an exact category ${\mathcal{C}}$ over an acyclic quiver $Q$. In [XZZ], the Auslander-Reiten translation formula was given for $Q$ an $\mathbb{A}_{n}$ type quiver with linear orientation and ${\mathcal{C}}=\Lambda\textendash\operatorname{mod}$ for some artin algebra $\Lambda$. In [GKKP3], the authors explicitly described the Auslander-Reiten translation for the monomorphism category of dualisable pro-species.
Especially, we are interested in the monomorphism categories ${\mathcal{S}}({\mathcal{C}})$ of Frobenius categories ${\mathcal{C}}$, which by [C1] are also Frobenius. As an application, when ${\mathcal{C}}$ is a stably $d$-Calabi-Yau Frobenius category, we calculate objects under the powers of Auslander-Reiten translation in the triangulated category $\overline{{\mathcal{S}}({\mathcal{C}})}$.
Theorem B.
Let ${\mathcal{C}}$ be a functorially finite stably $d$-Calabi-Yau Frobenius category of $\Lambda\textendash\operatorname{mod}$. Let $\langle 1\rangle$ be the suspension functor of the stable category $\overline{{\mathcal{S}}({\mathcal{C}})}$. The Auslander-Reiten translation $\tau_{\mathcal{S}}:\overline{{\mathcal{S}}({\mathcal{C}})}\to\overline{{\mathcal{S}}({\mathcal{C}})}$ satisfies $\tau_{\mathcal{S}}^{6}f\cong f\langle 6d-4\rangle$ for every object $f$ in $\overline{{\mathcal{S}}({\mathcal{C}})}$.
1.3. Outline of the paper
This paper is organized in the following way. In section 2, we recall the definitions of exact categories, stable categories and almost split sequences. In section 3, we recall the exact structures on morphism categories and characterize almost split sequences in morphism categories as well as their relation with almost split sequences in the functor categories. In section 4, we study the existence of almost split sequences in monomorphism categories and prove Theorem A(1). Section 5 is devoted to studying the minimal monomorphisms and epivalences. In section 6, we present the proof of Theorem A(2). In section 7, Auslander-Reiten translations of monomorphism categories of Frobenius categories are investigated. In section 8, we provide examples to further illustrate our results.
2. Almost split sequences in exact categories
2.1. Exact categories and (co)stable categories
First, we recall some fundamental notions of exact categories in the sense of Quillen [Q] and Keller [Ke, Appendix A]. One can also refer to [B] for a detailed survey.
An exact structure ${\mathcal{E}}$ on an additive category ${\mathcal{C}}$ is a fixed class of composable morphism pairs $A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C$, where $i$ is a kernel of $p$ and $p$ is a cokernel of $i$, satisfying certain axioms, see e.g. [B, Definition 2.1]. An exact category $({\mathcal{C}},{\mathcal{E}})$ is a pair of an additive category ${\mathcal{C}}$ and an exact structure ${\mathcal{E}}$ on it. An element $A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C$ of ${\mathcal{E}}$ is called a conflation or short exact sequence, where $i$ is called an inflaction or admissible monomorphism and $p$ is called a deflation or admissible epimorphism. Later on, when there is no confusion, we simply call ${\mathcal{C}}$ an exact category, omitting the exact structure ${\mathcal{E}}$ and call inflations (resp. deflations and conflations) monomorphisms (resp. epimorphisms and exact sequences). A conflation is denoted by $0\to A\stackrel{{\scriptstyle i}}{{\to}}B\stackrel{{\scriptstyle p}}{{\to}}C\to 0$.
Let $({\mathcal{C}},{\mathcal{E}})$ be an exact category. A full subcategory ${\mathcal{D}}\subseteq{\mathcal{C}}$ is called an exact subcategory of $({\mathcal{C}},{\mathcal{E}})$ if ${\mathcal{D}}$ is closed under extensions and equipped with an exact structure which contains $A\to B\to C\in{\mathcal{E}}$, where $A,B,C\in{\mathcal{D}}$.
In an exact category $({\mathcal{C}},{\mathcal{E}})$, an object $P$ is called projective if for any deflation $p:B\to C$, any morphism $u:P\to C$ can factor through $p$. An epimorphism $p:P(X)\to X$ is called a projective cover, if $P(X)$ is a projective object in ${\mathcal{C}}$ and any endomorphism $\phi\in\operatorname{End}_{\mathcal{C}}(P(x))$ satisfying $\phi=p\circ\phi$ is an automorphism. We say that an exact category $({\mathcal{C}},{\mathcal{E}})$ has enough projectives if for any object $C$ in ${\mathcal{C}}$, there exists a deflation $p:P\to C$ with $P$ a projective object. Notions with respect to injective objects and injective envelopes can be defined dually.
Remark 2.1.
In some literature, projective objects are also called $\operatorname{Ext}$-projective and exact subcategory are called fully exact subcategory
Let ${\mathcal{C}}$ be an exact category. Denote by $\underline{{\mathcal{C}}}$ the stable category, whose objects are the same as ${\mathcal{C}}$ and morphism sets are $\operatorname{Hom}_{\underline{{\mathcal{C}}}}(X,Y)=\operatorname{Hom}_{{\mathcal{C}}}(X,Y)/{\mathcal{P}}(X,Y)$, where ${\mathcal{P}}(X,Y)$ denotes the set of morphisms $X\to Y$ which can factor through a projective object. The costable category $\overline{\mathcal{C}}$ is defined dually. Morphisms in $\underline{{\mathcal{C}}}$ (resp. $\overline{\mathcal{C}}$) are usually denoted by $\underline{f}$ (resp.$\overline{f}$). Define $Q_{\mathcal{C}}:{\mathcal{C}}\to\underline{{\mathcal{C}}}$ the stable functor which sends morphisms $f\mapsto\underline{f}$. The costable functor $Q^{\mathcal{C}}:{\mathcal{C}}\to\overline{\mathcal{C}}$ is defined dually. A Frobenius category is an exact category $({\mathcal{C}},{\mathcal{E}})$ with enough projectives and enough injectives, moreover, the projectives coincide with the injectives.
Theorem 2.2.
[Hap1] Let ${\mathcal{C}}$ be a Frobenius category. Then $\underline{{\mathcal{C}}}=\overline{\mathcal{C}}$ is a triangulated category.
We are mostly interested in additive categories which are Krull-Schmidt. Recall that an additive category $\mathcal{C}$ is called Krull-Schmidt if any object $M\in\mathcal{C}$ has a unique (up to permutation) decomposition: $M=\mathop{\oplus}\limits_{i=1}^{n}M_{i}$ where $\operatorname{End}_{\mathcal{C}}(M_{i})$ is a local ring for all $1\leq i\leq n$.
The following result is well known (e.g. [KS, Proposition 1.2]), which guarantees the existence of projective covers and injective envelopes in Krull-Schmidt categories.
Proposition 2.3.
Let ${\mathcal{C}}$ be a Krull-Schmidt exact category with enough projective (resp. injective) objects. Then each object admits a projective cover (resp. injective envelope).
Denote by ${\mathcal{C}}_{\mathcal{P}}$ (resp. ${\mathcal{C}}_{\mathcal{I}}$) the full subcategory consisting of objects in ${\mathcal{C}}$ which do not contain any non-zero projective (resp. injective) summands. There is an equivalence $\underline{{\mathcal{C}}_{\mathcal{P}}}\cong\underline{{\mathcal{C}}}$ (resp. $\overline{{\mathcal{C}}_{\mathcal{I}}}\cong\overline{\mathcal{C}}$) induced by the inclusion ${\mathcal{C}}_{\mathcal{P}}\to{\mathcal{C}}$ (resp. ${\mathcal{C}}_{\mathcal{I}}\to{\mathcal{C}}$). The following result is easy to prove (cf.[ARS, IV. Exercise 7]).
Lemma 2.4.
Let ${\mathcal{C}}$ be a Krull-Schmidt exact category with enough projectives. A morphism $f:A\to B$ in ${\mathcal{C}}_{\mathcal{P}}$ is an isomorphism if and only if $\underline{f}$ is an isomorphism in $\underline{{\mathcal{C}}}$.
We introduce the notion of epivalences following [GKKP2, Definition 5.5].
Definition 2.5.
An additive functor $F:{\mathcal{C}}\to{\mathcal{D}}$ is said to be an epivalence, if it is full, dense and reflects isomorphisms.
An important fact about epivalences is that $F$ preserves and reflects indecomposable objects. Hence an epivalence $F:{\mathcal{C}}\to{\mathcal{D}}$ immediately yields a bijection between indecomposable objects in ${\mathcal{C}}$ and ${\mathcal{D}}$. Notice that the notion of epivalences was introduced in [Aus, Chapter II] by the name of “representation equivalence”. We prefer to use the name “epivalence”, because it emphasizes the fact that for an epivalence $F:{\mathcal{C}}\to{\mathcal{D}}$, ${\mathcal{C}}$ may contain more morphisms. Now we show an example of epivalence.
Lemma 2.6.
Let $\iota$ be the inclusion of ${\mathcal{C}}$ as a full subcategory of ${\mathcal{D}}$. The restriction $Q_{{\mathcal{C}}}|_{{\mathcal{C}}_{{\mathcal{P}}}}=Q_{\mathcal{C}}\circ\iota$ is an epivalence.
Proof.
It is easy to see $Q_{\mathcal{C}}\circ\iota$ is full and dense. It reflects isomorphisms due to Lemma 2.4. ∎
Corollary 2.7.
The functor $Q_{\mathcal{C}}|_{{\mathcal{C}}_{{\mathcal{P}}}}$ induces a bijection between indecomposable objects in ${\mathcal{C}}_{\mathcal{P}}$ and $\underline{{\mathcal{C}}}$.
2.2. Almost split sequences
Following [INP, Definition 3.4], we introduce the Auslander-Reiten-Serre duality for exact categories. An additive category ${\mathcal{C}}$ is $k$-linear if its hom-sets are $k$-modules. Denote by $D=\operatorname{Hom}_{k}(-,k)$ the duality functor.
Definition 2.8.
Let ${\mathcal{C}}$ be a $k$-linear exact category.
(1) A right Auslander-Reiten-Serre (ARS) duality is a pair $(\tau,\eta)$ of an additive functor $\tau:\underline{{\mathcal{C}}}\to\overline{\mathcal{C}}$ and a binatural isomorphism $\eta_{A,B}:\operatorname{Hom}_{\underline{{\mathcal{C}}}}(A,B)\cong D\operatorname{Ext}^{1}_{\mathcal{C}}(B,\tau A)$ for any $A,B\in{\mathcal{C}}$
(2) If moreover $\tau$ is an equivalence, we say that $(\tau,\eta)$ is an Auslander-Reiten-Serre (ARS) duality.
Definition 2.9.
Let ${\mathcal{C}}$ be a Krull-Schmidt exact category. We call $f:A\to B$ in ${\mathcal{C}}$ a source map or left almost split, if it is not a section and each $u:A\to B^{\prime}$ which is not a section factors through $f$. Dually, we call $g:B\to C$ a sink map or right almost split, if it is not a retraction and each $v:B^{\prime}\to C$ which is not a retraction factors through $g$.
An almost split sequence in ${\mathcal{C}}$ is a conflation $A\stackrel{{\scriptstyle f}}{{\to}}B\stackrel{{\scriptstyle g}}{{\to}}C$ such that $f$ is a source map and $g$ is a sink map.
We say ${\mathcal{C}}$ has almost split sequences if it satisfies:
  1. (1)
    Each indecomposable object $A$ admits a source map $A\to B$ and a sink map $B\to A$.
  2. (2)
    For each indecomposable non-projective object $C$, there is an almost split sequence $A\to B\to C$.
  3. (3)
    For each indecomposable non-injective object $A$, there is an almost split sequence $A\to B\to C$.
Let ${\mathcal{C}}$ be a Hom-finite $k$-linear Krull-Schmidt category. Denote by ${\mathcal{C}}\textendash\operatorname{Mod}$ the category of additive contravariant functors from ${\mathcal{C}}$ to abelian groups and ${\mathcal{C}}\textendash\operatorname{mod}$ the subcategory consisting of finitely presented functors. Then ${\mathcal{C}}$ is called a dualizing $k$-variety [AR74] if the $k$-dual functors $D:{\mathcal{C}}\textendash\operatorname{Mod}\to{\mathcal{C}}^{\rm op}\textendash\operatorname{Mod}$ and $D:{\mathcal{C}}^{\rm op}\textendash\operatorname{Mod}\to{\mathcal{C}}\textendash\operatorname{Mod}$ given by $D(F)(C)=D(FC)$ for every object $C$ of ${\mathcal{C}}$ and $F\in{\mathcal{C}}\textendash\operatorname{Mod}$ or ${\mathcal{C}}^{\rm op}\textendash\operatorname{Mod}$ induce dualities
$D:{\mathcal{C}}\textendash\operatorname{mod}\to{\mathcal{C}}^{\rm op}\textendash\operatorname{mod}\ \mbox{ and }\ D:{\mathcal{C}}^{\rm op}\textendash\operatorname{mod}\to{\mathcal{C}}\textendash\operatorname{mod}.$
In this case, it turns out that ${\mathcal{C}}\textendash\operatorname{mod}$ is an abelian subcategory of ${\mathcal{C}}\textendash\operatorname{Mod}$ that admits enough projective and enough injective objects [AR74, Theorem 2.4]. We also note that if ${\mathcal{C}}$ is a dualizing $k$-variety then so is the functor category ${\mathcal{C}}\textendash\operatorname{mod}$ ([AR74, Proposition 2.6]). In particular, ${\mathcal{C}}\textendash\operatorname{mod}$ is also Krull-Schmidt. Furthermore, any functorially finite subcategory of a dualizing $k$-variety is a dualizing $k$-variety by [AS, Theorem 2.3].
Theorem 2.10.
[INP, Corollary 4.5] Let ${\mathcal{C}}$ be a $k$-linear $\operatorname{Ext}$-finite Krull-Schmidt exact category with enough projectives and injectives. Then the following conditions are equivalent.
  1. (1)
    ${\mathcal{C}}$ has almost split sequences.
  2. (2)
    ${\mathcal{C}}$ has an Auslander-Reiten-Serre duality.
  3. (3)
    The stable category $\underline{{\mathcal{C}}}$ is a dualizing $k$-variety.
  4. (4)
    The stable category $\overline{\mathcal{C}}$ is a dualizing $k$-variety.
3. Morphism categories and functor categories
Let ${\mathcal{C}}$ be an additive category. The objects in the morphism category ${\mathcal{H}}({\mathcal{C}})$ are morphisms $f:X_{1}\to X_{2}$ in $\mathcal{C}$ also denoted by $\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)_{f}$. A morphism in ${\mathcal{H}}({\mathcal{C}})$
$\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right):\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)_{f}\to{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}$
is a pair of ${\mathcal{C}}$-morphisms $\phi_{1}$, $\phi_{2}$ such that $\phi_{2}f=g\phi_{1}$.
3.1. Exact structures on morphism categories
Definition 3.1 (The cw-exact structure).
Let ${\mathcal{E}}^{\rm cw}$ be the class of all pairs of composable morphisms
${\delta:}$${{\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}}$
such that the induced composable morphisms $X_{i}\stackrel{{\scriptstyle\phi_{i}}}{{\rightarrow}}Y_{i}\stackrel{{\scriptstyle\psi_{i}}}{{\rightarrow}}Z_{i}$ split in ${\mathcal{C}}$ for $i=1,2$. We call the elements in ${\mathcal{E}}^{\rm cw}$ cw-exact sequences.
It is worth mentioning the following easy and useful observation due to [HE].
Remark 3.2.
Any cw-exact sequence
${\delta:}$${{\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}}$
is isomorphism to
${\delta^{\prime}:}$${{\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\left[\begin{smallmatrix}1\\ 0\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}1\\ 0\end{smallmatrix}\right]\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}X_{1}\oplus Z_{1}\\ X_{2}\oplus Z_{2}\end{smallmatrix}\right)}_{g^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}0&1\end{smallmatrix}\right]\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}}$
where $g^{\prime}=\begin{bmatrix}f&q\\ 0&h\end{bmatrix}$ with $q:Z_{1}\to X_{2}$ a possibly non-zero morphism in ${\mathcal{C}}$.
Recall that an additive category $\mathcal{C}$ is called idempotent-complete if every idempotent admits a kernel. In particular, if $\mathcal{C}$ is an exact subcategory of $\Lambda\textendash\operatorname{mod}$, $\mathcal{C}$ is idempotent-complete if and only if it is closed under direct summands.
Proposition 3.3.
[HE, Proposition 3.2] Assume $\mathcal{C}$ is idempotent-complete and admits pseudo-kernels. Then ${\mathcal{E}}^{\rm cw}$ defines an exact structure on the additive category ${{\mathcal{H}}}({\mathcal{C}})$.
When ${\mathcal{C}}$ satisfies the above proposition, we denote the exact category $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm cw})$ by ${{\mathcal{H}}}^{\rm cw}({\mathcal{C}})$.
If ${\mathcal{C}}$ is moreover an exact category, then ${\mathcal{H}}({\mathcal{C}})$ enjoys a richer exact structure, which is induced naturally from the exact structure of ${\mathcal{C}}$.
Definition 3.4 (The deg-exact structure).
Let ${\mathcal{E}}^{\rm deg}$ be the class of all pairs of composable morphisms
${\delta:}$${{\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}}$
such that the induced composable morphisms $X_{i}\stackrel{{\scriptstyle\phi_{i}}}{{\rightarrow}}Y_{i}\stackrel{{\scriptstyle\psi_{i}}}{{\rightarrow}}Z_{i}$ exact in ${\mathcal{C}}$ for $i=1,2$.
When ${\mathcal{C}}$ is exact, there are two exact structures on ${\mathcal{H}}({\mathcal{C}})$, namely ${\mathcal{E}}^{\rm cw}$ and ${\mathcal{E}}^{\rm deg}$. It is clear that ${\mathcal{E}}^{\rm cw}\subseteq{\mathcal{E}}^{\rm deg}$. So any cw-exact sequence is still exact in $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$. Intuitively, ${\mathcal{E}}^{\rm deg}$ is the more natural exact structure because when ${\mathcal{C}}\subseteq\Lambda\textendash\operatorname{mod}$, $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ is an exact subcategory of ${\mathcal{H}}(\Lambda)\cong T_{2}(\Lambda)\textendash\operatorname{mod}$. Hence, in the following, when we refer to the exact category ${{\mathcal{H}}}({\mathcal{C}})$ without emphasizing the exact structure, we mean $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$. Although the exact structure ${\mathcal{E}}^{\rm cw}$ is much smaller than ${\mathcal{E}}^{\rm deg}$ in general, it is surprising that when $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ has almost split sequences, most of the almost split sequences are cw-exact. Hence, when it comes to characterizing almost split sequences, ${\mathcal{H}}^{\rm cw}({\mathcal{C}})$ becomes a convenient tool.
3.2. Functors on morphisms categories
Let $F:{\mathcal{C}}\to{\mathcal{D}}$ be an additive covariant functor between additive categories ${\mathcal{C}}$ and ${\mathcal{D}}$. Then $F$ naturally induces an additive functor ${\mathcal{H}}(F):{\mathcal{H}}({\mathcal{C}})\to{\mathcal{H}}({\mathcal{D}})$, which sends $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)_{f}$ to $\left(\begin{smallmatrix}FA\\ FB\end{smallmatrix}\right)_{Ff}$ and sends morphisms $(\alpha,\beta):f\to g$ to $(F\alpha,F\beta):Ff\to Fg$.
Proposition 3.5.
Let $F:{\mathcal{C}}\to{\mathcal{D}}$ be an additive functor. If $F$ is an equivalence, then so is ${\mathcal{H}}(F)$.
Proof.
It suffices to prove ${\mathcal{H}}(F)$ is dense and fully-faithful. For the density of ${\mathcal{H}}(F)$, let $\left(\begin{smallmatrix}A^{\prime}_{1}\\ A^{\prime}_{2}\end{smallmatrix}\right)_{g}\in{\mathcal{H}}({\mathcal{D}})$. By the density of $F$, there are $A_{1},A_{2}\in{\mathcal{C}}$, such that $F(A_{1})\cong A^{\prime}_{1},F(A_{2})\cong A^{\prime}_{2}$. Since $F$ is full, there is a homomorphism $f:A_{1}\to A_{2}$ such that $F(f)\cong g.$ So ${\mathcal{H}}(F)$ is dense. For the fullness of ${\mathcal{H}}(F)$, let $\left(\begin{smallmatrix}A_{1}\\ A_{2}\end{smallmatrix}\right)_{f}$ and $\left(\begin{smallmatrix}B_{1}\\ B_{2}\end{smallmatrix}\right)_{g}$ be objects in ${\mathcal{C}}$ and $(\alpha^{\prime},\beta^{\prime}):Ff\to Fg$ be a homomorphism in ${\mathcal{H}}({\mathcal{D}})$, i.e. $Fg\alpha^{\prime}=\beta^{\prime}Ff$. By the fullness of $F$, there are homomorphisms $\alpha:A_{1}\to A_{2}$ and $\beta:B_{1}\to B_{2}$ such that $F\alpha=\alpha^{\prime}$ and $F\beta=\beta^{\prime}$. Hence $FgF\alpha=F\beta Ff$. Since $F$ is a faithful functor, we have $g\alpha=\beta f$. That is to say, $(\alpha,\beta):f\to g$ is a homomophism in ${\mathcal{H}}({\mathcal{C}})$ such that $F(\alpha,\beta)=(\alpha^{\prime},\beta^{\prime})$. Hence ${\mathcal{H}}(F)$ is full. It is easy to see that the faithfulness of ${\mathcal{H}}(F)$ is guaranteed by the faithfulness of $F$. This completes the proof. ∎
Let ${\mathcal{C}}$ be an exact category. Define the monomorphism category $\mathcal{S}({\mathcal{C}})$ (resp. epimorphism category $\mathcal{F}({\mathcal{C}})$) as the full subcategory of $(\mathcal{H}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$, consisting of objects which are monomorphisms (resp. epimorphisms).
There are a couple of functors:
$\operatorname{\bf Ker}:{\mathcal{F}}({\mathcal{C}})\to{\mathcal{S}}({\mathcal{C}}),$ $\operatorname{\bf Cok}:{\mathcal{S}}({\mathcal{C}})\to{\mathcal{F}}({\mathcal{C}})$
$\left(\begin{matrix}A\\ B\end{matrix}\right)_{f}\mapsto\left(\begin{matrix}\ker f\\ A\end{matrix}\right)_{i},$ $\left(\begin{matrix}A\\ B\end{matrix}\right)_{f}\mapsto\left(\begin{matrix}B\\ \operatorname{Coker}f\end{matrix}\right)_{p}$
where $i$ is the canonical embedding and $p$ is the canonical projection. The following result is usually referred to as the Ringel-Schmidmeier-Simson (RSS) equivalence:
Proposition 3.6.
Additive functors $\operatorname{\bf Ker}$ and $\operatorname{\bf Cok}$ form a pair of inverse equivalences.
As immediate consequences of the RSS equivalence. We will see that $\operatorname{\bf Ker}$ and $\operatorname{\bf Cok}$ preserve indecomposable objects, projectives, injectives and almost split sequences.
3.3. Almost split sequences in morphism categories
The following result is due to [HE, Proposition 4.6], which asserts the existence of almost split sequences in the morphism categories.
Proposition 3.7.
Let ${\mathcal{C}}$ be an extension-closed $k$-dualizing subvariety of $\Lambda\textendash\operatorname{mod}$. Then the canonical exact category $(\mathcal{H}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ admits almost split sequences.
Definition 3.8.
An indecomposable object in ${\mathcal{H}}({\mathcal{C}})$ is said to be of type $(a)$ (resp. $(b)$, or $(c)$) provided it is isomorphic to $(0\rightarrow M)$ (resp. $(M\stackrel{{\scriptstyle 1}}{{\rightarrow}}M)$, or $(M\rightarrow 0)$) for some $\Lambda$-module $M$. Furthermore, an indecomposable object is said to be of type $(d)$ if it is isomorphic to $(P\stackrel{{\scriptstyle f}}{{\rightarrow}}Q)$ where $P,Q$ are projective $\Lambda$-modules.
The following characterization of almost split sequences can be obtained from the definitions.
Lemma 3.9.
[HE, Lemma 4.5] Assume $\delta:0\rightarrow A\stackrel{{\scriptstyle f}}{{\rightarrow}}B\stackrel{{\scriptstyle g}}{{\rightarrow}}C\rightarrow 0$ is an almost split sequence in ${\mathcal{C}}$. Then
  • $(1)$
    The almost split sequence in ${{\mathcal{H}}}({\mathcal{C}})$ ending at $(0\rightarrow C)$ is of the form
[…]
  • $(2)$
    The almost split sequence in ${{\mathcal{H}}}({\mathcal{C}})$ ending at $(C\stackrel{{\scriptstyle 1}}{{\rightarrow}}C)$ is of the form
[…]
The following corollary is an immediate consequence of the arguments above.
Corollary 3.10.
Let ${\mathcal{C}}$ be an extension-closed $k$-dualizing subvariety of $\Lambda\textendash\operatorname{mod}$ and let
[…]